OpenStudy (anonymous):

Solve by doubling both side: cos x - sin x = sqrt 2/2

OpenStudy (anonymous):

that's (sqrt 2)/2

OpenStudy (mathstudent55):

2(cos x - sin x) = sqrt(2) ( 2(cos x - sin x) )^2 = 2 4(cos^2 x - 2sinxcosx + sin^2) = 2 sin^2x + cos^2x - 2sinxcosx = 1/2 1 - 2sinxcosx = 1/2 -sinxcosx = -1/2 sinxcosx = 1/2

OpenStudy (mathstudent55):

2sinxcosx = 1 sin2x = 1

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

would 2sinxcos and 2cosxsinx the same thing?

OpenStudy (anonymous):

just switched around a bit

OpenStudy (mathstudent55):

Yes. 2sinxcosx = sin2x. This is one of the double angle identities. Now that you have sin2x = 1, that will give you 2x easily. Then divide the result by 2 to get x. Remember that if you are solving this in an interval, such as [0, 2pi], you have one answer. If you are solving it in general, then you have to add a term with n (any integer), to account for the periodic repetition of the solution.